OpenStudy (anonymous):

The line of symmetry for the quadratic equation y = ax^2 - 8x - 3 is x = 2. What is the value of "a"? 2?

ganeshie8 (ganeshie8):

hint- (2,0) lies on the quadratic equation

OpenStudy (anonymous):

-2?

OpenStudy (anonymous):

dont think 19/3 is the answer its either 2, -2 and -1

ganeshie8 (ganeshie8):

nope pls try again buddy... :) go step by step.... .

ganeshie8 (ganeshie8):

plugin y=0, x = 2 0 = a(2^2) - 8(2) - 3 0 = 4a-16-3 0 = 4a-19 19 = 4a 19/4 = a

mathslover (mathslover):

check ur process again @ineedhelpplease57

OpenStudy (anonymous):

ok..

mathslover (mathslover):

did u get the mistake?

OpenStudy (anonymous):

yes